how do you calculate the angles of a kite when 2 angles(B and D) are equal and both equal 114 degrees and one other angle (A)is twice the measure of angle C

Found 2 solutions by scott8148, Edwin McCravy:
Answer by scott8148(6628)

the sum of the angles of a quadrilateral (like a kite) is 360º

3C = 360 - 2(114)

Answer by Edwin McCravy(20086)



             A + B + C + D = 360°  

Substitute B = 114°, D = 114°, and A = 2C

      2C + 114° + C + 114° = 360°

                 3C + 228° = 360°

                        3C = 132°

                         C = 44° 

So angle A = 2C = 2(44°) = 88°

Angle B = 114°

Angle C = 44°

Angle D = 114°

Edwin